Respected tutor iam in hurry and want answer plz help me.Thankyou.The question is:Find two consecutive natural numbers,such that one third of the smaller numbers is one more than one fourth of the greater.
Answer by nerdybill(5410)

Find two consecutive natural numbers,such that one third of the smaller numbers is one more than one fourth of the greater.
.
Let x = first of two consecutive numbers (smaller)
then
x+1 = second of two consecutive numbers (larger)
.
(1/3)x = (1/4)(x+1)
Multiplying both sides by 12 (common denominator):
4x = 3(x+1)
4x = 3x+3
x = 3 (smaller number)
.
Larger number:
x+1 = 3+1 = 4
.
The two consecutive numbers are: 3 and 4